OpenStudy (anonymous):

Help! ^_^ A business consultant earns a flat fee for his work as well as an hously fee. He charges his clients at a rate of $75 per hour. In one week, he billed a client $1,800 for 20 hours worked. Which of the following equations describes the pay for the consultant in terms of hours worked? a. y=20x-300 b. y=20x+300 c. y=75x-300 d. y=75x+300

jimthompson5910 (jim_thompson5910):

"He charges his clients at a rate of $75 per hour" so m = 75 which means y = mx+b becomes y = 75x+b

jimthompson5910 (jim_thompson5910):

"he billed a client $1,800 for 20 hours worked" tells us that when x = 20, y = 1800 So y = 75x+b turns into 1800 = 75*20+b Solve that for b Then use m = 75 and that value of b to find the equation

OpenStudy (anonymous):

I like have no idea how you do this ! :/

jimthompson5910 (jim_thompson5910):

which part are you stuck on?